OpenStudy (anonymous):

Use the triangle at the right. Find the length of the missing side. 1. a = 16, b = 63 2. b = 2.1, c = 2.9 i need help

OpenStudy (e.mccormick):

OK. So what triangle?

OpenStudy (e.mccormick):

If it is a right triangle, you can use the Pythagorean Theorem. But for other things, you may need trig.

OpenStudy (anonymous):

i need help with all of this i dont understand any of it

OpenStudy (e.mccormick):

Ah, right triangles! Pythagorean Theorem it is!  If this is basically any right triangle, then \(x^2+y^2=h^2\) h would be the hypotenuse. Familliar with that one?

OpenStudy (anonymous):

yes

OpenStudy (e.mccormick):

Great! Well, in your problem, they are using a, b, and c. Which one is the hypotenuse? Once you know that, it is just a matter of writing it properly and solving for the unknown.

OpenStudy (anonymous):

the hypotenuse is c

OpenStudy (e.mccormick):

OK. So the formula becomes: \(a^2+b^2=c^2\) So, replace a, b, and c with what you know. In #1 it is a = 16, b = 63, and solve for c. In #2 it is b = 2.1, c = 2.9, solve fora.

OpenStudy (anonymous):

so ifon number 1 a=16 and b+63 then c the hypotenuse =79

OpenStudy (anonymous):

and if b= 2.1 and c= 2.9 then a=.8

OpenStudy (e.mccormick):

You need to use the squares. Let me show you the first one. \(a^2+b^2=c^2 \implies (16)^2+(63)^2=c^2 \implies 256+ 3969 = c^2\) See where that is going?

OpenStudy (e.mccormick):

So I have \(4225=c^2\) but what I need is c. So then I take the root. \(\sqrt{4225}=c\) If I find the root, I am done.

OpenStudy (e.mccormick):

So \(c=65\). Do you see why the squares make such a huge difference in the answer?

OpenStudy (anonymous):

yea

OpenStudy (e.mccormick):

OK. So do the second one again, but remember to square, do the addition or subtraction in that case, then find the root.

OpenStudy (e.mccormick):

Ooh. Nice number to get a root of in that one. Should be an easy answer once you do the squares properly.

OpenStudy (anonymous):

ok im lost im problem 2 ivde tried and messed up

OpenStudy (e.mccormick):

What did you try?

OpenStudy (e.mccormick):

Had some connection problems... Look at this: \(a^2+(2.1)^2=(2.9)^2\) You need to square the numbers. Then, whatever you get for the replaced b needs to be moved over to the other side, which means subtracting it. Finally, you will do the root of both sides to turn \(b^2\) into b and whatever number is on the right into the answer. I have to go now, but I hope that is enough to get you through. The next question, with the ladder, it the same basic thing. Just need to pick out what to put into the formula.